# A Scalable Unsupervised Framework for multi-aspect labeling of Multilingual and Multi-Domain Review Data

### v1

@@ 5.5 Comparison of Multi-Label Classification and LLM Results
To evaluate the performance of MUSCAD, we conducted comparative experiments with large language models (LLMs). This was done to assess the feasibility of LLM-based labeling and to verify how consistently and accurately fine-tuned models can perform labeling. We extracted 1,000 samples per domain for evaluation. Table 6 and Table 7 compare the results of multi-aspect labeling performed by fine-tuned models and LLMs (GPT-3.5-turbo and GPT-4o-mini) for the beauty, hotel, and food domains, respectively. The LLM-based labeling was conducted using prompts under zero-shot and few-shot settings, and the prompts are provided in D.
@@
 Performance  Fine-tuned models performed labeling based on data trained for specific domains, and as a result, generally achieved higher Micro-F1 and Macro-F1 scores. In particular, KR-BERT and XLM-RoBERTa showed the best performance in the hotel and food domains, while BERT-large and RoBERTa performed relatively well in the beauty domain. On the other hand, the labeling results using LLMs were generally inferior to those of the fine-tuned models. Although the LLMs demonstrated the ability to interpret sentence meaning and assign appropriate aspects, issues were observed such as assigning incorrect labels or producing inconsistent results for the same sentence (see Table 8).
@@
Two men traveled, and perhaps because it was off-season, there were hardly any people; it was quiet and offered good value for the price. Satisfaction Purpose Satisfaction Location
The tangsuyuk is also tasty, but since it’s located on the first floor of a building where many companies are situated, there’s a really long wait at lunchtime. Food Satisfaction Waiting Time Taste Location Waiting Time
I returned this item because it had terrible coverage, similar to using a skin tone primer Color Satisfaction () Satisfaction
Table 8: Comparison of Aspect Category Labeling by MUSCAD and an LLM-Based Model. A green check () indicates a correctly identified aspect, whereas a red cross () denotes an incorrect or missing aspect.
@@
In conclusion, while LLM-based multi-aspect labeling can be useful for quickly labeling small datasets, fine-tuned models prove to be a more practical solution for large-scale data processing. They demonstrate higher reproducibility and more consistent performance compared to LLMs, and are more cost-effective, making them better suited for real-world applications.
@@ 5.7 Summary
As demonstrated through various experiments in Section 5, the multi-aspect labels generated by the proposed framework are of sufficient quality for use in model training. The multi-label classification experiments using various fine-tuned models yielded high Micro-F1 and Macro-F1 scores, indicating that the automatically labeled data is suitable for supervised learning. Furthermore, we compared the performance of fine-tuned models with few-shot labeling using GPT-3.5-turbo and GPT-4o-mini to evaluate the capabilities of LLMs in multi-aspect labeling. The results showed that while LLMs were able to assign appropriate labels in some cases, their overall performance was inferior to that of the fine-tuned models. LLMs also exhibited issues with reproducibility and incurred higher costs. In addition, Human Evaluation confirmed that the automatic labels produced by MUSCAD achieved a level of reliability and consistency comparable to manual labeling by experts, and were more effective in reducing subjectivity among evaluators. These findings suggest that the proposed framework offers a more practical solution for applying multi-aspect labeling to large-scale review data.
@@ 6.1 Conclusion
In this study, we proposed MUSCAD, an unsupervised multi-aspect labeling framework designed to automatically assign aspect categories to multilingual and multi-domain review data. While prior research has largely relied on supervised learning approaches that require large amounts of manually labeled data, our framework integrates K-means–based aspect candidate generation, contextual representation learning using Multi-Head Attention, and classification optimization via Max-Margin Loss, enabling multi-aspect labeling without predefined vocabularies. Experimental results demonstrated that fine-tuned classification models trained on the resulting multi-aspect labeling dataset achieved high F1-scores, validating the reliability of the automatically generated labels. Compared to large language models (LLMs), MUSCAD provided more consistent and reproducible labeling outcomes. In addition, Human Evaluation confirmed that the quality of the automatically generated labels was comparable to that of manual labeling by experts. Finally, the MUSCAD  framework has shown scalability and practical applicability in real-world environments through large-scale automatic labeling across multiple languages and domains.
@@ 6.2 Future Work
This study proposed and evaluated the performance of the unsupervised MUSCAD framework across diverse domains. However, several directions remain for future improvements and extensions. First, we aim to build a fully automated labeling system using AI agents. Currently, partial manual intervention is required for matching aspect categories. In future work, we plan to eliminate manual processes entirely by incorporating AI agents and establishing an automated feedback loop, ultimately enabling fully autonomous labeling. Second, developing aspect-aware review summarization techniques is another key direction. While this study focuses on identifying the aspects discussed in each review, it does not provide functionality for summarizing the core content of individual reviews. As a next step, we aim to build a model that generates aspect-level summaries using the labeled data. For example, in hotel reviews, aspects such as “Room”, “Location” and “Service” can be automatically extracted, and the content related to each aspect can be summarized and presented to users. This would help enhance user experience and improve the utility of review data. Building upon these directions, we intend to further refine MUSCAD and develop an automated labeling and summarization system that can be applied in real-world settings. By integrating a fully automated labeling pipeline based on AI agents with aspect-aware review summarization, we envision the creation of a more efficient and scalable NLP-based review analysis system.
@@ Appendix C Multi-Aspect Category Labeling Results
This appendix presents the prompts used for multi-aspect labeling with LLMs in the comparative experiments described in Section 5.5. In this study, we designed appropriate prompts to evaluate whether LLMs can assign correct aspect categories based on review sentences under zero-shot and few-shot settings. The prompts are shown in Fig. D.1, Fig. D.2, Fig. D.3, Fig. D.4, and Fig. D.5.
@@
Appendix D Prompt
This appendix presents the prompts used for multi-aspect labeling with LLMs in the comparative experiment described in Section 5.5. In this study, we designed appropriate prompts to evaluate whether LLMs can assign correct aspect categories to review sentences under zero-shot and few-shot settings. The prompts are shown in Fig. D.1, Fig. D.2, Fig. D.3, Fig. D.4, and Fig. D.5.

### v2

@@ 5.5 Comparison of Multi-Label Classification and LLM Results
To evaluate the performance of MUSCAD, we conducted comparative experiments with large language models (LLMs). Recent NLP research has emphasized the contrast between domain-specialized fine-tuned models and general-purpose LLMs, showing that fine-tuned models often achieve higher robustness and task-specific accuracy [40]. Building on this line of research, our experiments were designed to assess whether this trend also holds in the context of multi-aspect labeling and to empirically validate the performance of the proposed MUSCAD framework against LLM-based labeling approaches. Specifically, we compared models fine-tuned on the automatically labeled datasets generated by the proposed MUSCAD framework with prompted LLMs, analyzing the consistency and accuracy of each approach across multiple domains. We extracted 1,000 samples per domain for evaluation. Table 6 and Table 7 compare the results of multi-aspect labeling performed by fine-tuned models and LLMs (GPT-3.5-turbo and GPT-4o-mini) for the beauty, hotel, and food domains, respectively. The LLM-based labeling was conducted using prompts under zero-shot and few-shot settings, and the prompts are provided in E.
@@
Performance  Fine-tuned models, which perform labeling based on data trained for specific domains, generally achieved higher Micro-F1 and Macro-F1 scores. In particular, KR-BERT and XLM-RoBERTa showed the best performance in the hotel and food domains, while BERT-large and RoBERTa performed relatively well in the beauty domain. On the other hand, the labeling results using LLMs were generally inferior to those of the fine-tuned models. Although the LLMs demonstrated the ability to interpret sentence meaning and assign appropriate aspects, issues were observed such as assigning incorrect labels or producing inconsistent results for the same sentence (see Table 8).
@@
Two men traveled, and perhaps because it was off-season, there were hardly any people; it was quiet and offered good value for the price. Satisfaction ✓ Purpose ✓ Satisfaction ✓ Location
The tangsuyuk is also tasty, but since it’s located on the first floor of a building where many companies are situated, there’s a really long wait at lunchtime. Food ✓ Satisfaction ✓ Waiting Time ✓ Taste Location Waiting Time ✓
I returned this item because it had terrible coverage, similar to using a skin tone primer Color ✓ Satisfaction ✓ () Satisfaction ✓
Table 8: Comparison of Aspect Category Labeling by MUSCAD and an LLM-Based Model. A green check (✓) indicates a correctly identified aspect, whereas a red cross () denotes an incorrect or missing aspect.
@@
In conclusion, while LLM-based multi-aspect labeling can be useful for quickly labeling small datasets, fine-tuned models prove to be a more practical solution for large-scale data processing. Moreover, these results are consistent with recent studies that highlight the instruction-following limitations of large language models [39]. That work points out the fundamental limitation of LLMs in strictly adhering to complex task instructions and suggests that additional fine-tuning approaches are required to mitigate this issue. Our study also observed similar limitations while LLMs exhibit strong contextual understanding, they still show weaknesses in terms of output stability and reproducibility. They demonstrate higher reproducibility and more consistent performance compared to LLMs, and are more cost-effective, making them better suited for real-world applications.
@@ 5.6 Human Evaluation
5.7 Comparison with Unsupervised Aspect Extraction Methods
Domain Model Type Model NPMI UMass Diversity EC Rank Avg
Hotel Probabilistic-based LDA 0.026 -3.833 1.000 5.7
BTM -0.337 -18.132 0.143 9.7
NMF 0.038 -3.770 0.671 7.0
Hybrid CTM 0.195 -0.229 0.871 0.499 3.3
Embedding-based BERTopic 0.173 -0.480 0.692 0.550 4.3
Top2Vec 0.025 -0.683 0.094 0.502 7.3
KeyBERT 0.087 -0.382 0.707 0.266 5.3
SimCSE -0.018 -0.534 0.343 0.520 6.5
ABAE 0.242 -0.266 0.863 0.486 3.8
MUSCAD 0.281 -0.220 1.000 0.568 1.0
Food Probabilistic-based LDA 0.021 -4.302 1.000 5.0
BTM -0.459 -20.227 0.111 10.0
NMF 0.017 -3.994 0.656 6.7
Hybrid CTM 0.222 -0.279 0.715 0.500 4.3
Embedding-based BERTopic 0.189 -0.409 0.528 0.561 4.3
Top2Vec 0.021 -0.592 0.098 0.510 6.0
KeyBERT 0.078 -0.630 0.705 0.508 5.0
SimCSE -0.015 -0.763 0.310 0.527 7.3
ABAE 0.334 -0.052 0.991 0.491 3.3
MUSCAD 0.384 -0.050 1.000 0.592 1.0
Beauty Probabilistic-based LDA 0.017 -4.416 1.000 5.7
BTM -0.493 -20.837 0.100 10.0
NMF 0.018 -3.923 0.600 7.0
Hybrid CTM 0.267 -0.123 0.939 0.281 3.8
Embedding-based BERTopic 0.146 -0.357 0.769 0.292 4.8
Top2Vec 0.102 -0.918 0.161 0.325 6.0
KeyBERT 0.170 -0.521 0.856 0.279 5.0
SimCSE -0.008 -0.667 0.444 0.332 7.0
ABAE 0.313 -0.016 1.000 0.274 3.0
MUSCAD 0.384 -0.008 1.000 0.360 1.0
Table 10: Comparison of Topic Modeling Methods across Domains
To objectively evaluate the quality of automatically extracted Aspect Terms, it is essential to employ evaluation metrics that can measure both the semantic coherence among Aspect Terms and the distinctiveness between different Aspect Categories, even in the absence of labeled data. [42][43] proposed coherence- and diversity-based evaluation metrics that have been widely adopted as quantitative measures showing strong correlations with human judgments of topic interpretability. Building upon these approaches, we comprehensively evaluate the performance of MUSCAD using four complementary metrics: NPMI, UMass, Diversity, and Embedding Coherence. NPMI and UMass assess the semantic and document-level coherence among top-ranked Aspect Terms, Diversity measures the redundancy across Aspect Terms, and Embedding Coherence evaluates the degree of semantic alignment in the embedding space. Together, these four metrics provide complementary perspectives that enable a holistic evaluation of how effectively MUSCAD learns coherent and well-separated Aspect representations in an unsupervised setting.
As introduced above, the topic coherence metrics, NPMI and UMass, are used to quantitatively evaluate the semantic consistency among the top-ranked words within each Aspect. In this study, both NPMI and UMass metrics were employed to achieve a balanced evaluation of semantic relatedness and document-level co-occurrence consistency. The Pointwise Mutual Information (PMI) measures the association strength between two words based on the ratio of their joint probability $P(w_{i},w_{j})$ to the product of their individual probabilities $P(w_{i})$ and $P(w_{j})$ . A small smoothing constant $\epsilon=1$ is applied to prevent zero probabilities during the logarithmic operation.
$\text{PMI}(w_{i},w_{j})=\log_{2}\left(\frac{P(w_{i},w_{j})+\epsilon}{P(w_{i})P(w_{j})}\right)$ (7)
The Normalized Pointwise Mutual Information (NPMI) is widely used because it shows a strong correlation with human judgments of topic interpretability [42]. NPMI extends PMI by normalizing its value to the range [1,1][-1,1], providing a more interpretable scale for assessing the strength of word associations. The NPMI between two words $w_{i}$ and $w_{j}$ is defined as follows:
$\text{NPMI}(w_{i},w_{j})=\frac{\log\frac{P(w_{i},w_{j})+\epsilon}{P(w_{i})P(w_{j})}}{-\log(P(w_{i},w_{j})+\epsilon)}$ (8)
Here, $P(w_{i},w_{j})$ denotes the probability that the two words co-occur within a sliding window, and $\epsilon$ is a smoothing constant used to avoid zero probabilities. A higher NPMI value indicates stronger semantic associations among Aspect Terms, implying greater intra-aspect cohesion.
UMass Coherence evaluates topic quality based on the co-occurrence of top-ranked words within the same document [43]. While NPMI relies on mutual information, UMass measures the average pairwise log conditional probability between words. It is defined as follows:
$\text{UMass}(w_{i},w_{j})=\log\frac{D(w_{i},w_{j})+\epsilon}{D(w_{j})}$ (9)
Here, $D(w_{i},w_{j})$ denotes the number of documents in which the two words co-occur, and $D(w_{j})$ represents the number of documents containing the word $w_{j}$. The small constant $\epsilon$ is added to avoid division by zero. A UMass value closer to zero indicates higher topic coherence.
Diversity measures the degree of redundancy among aspect categories. It is computed as the ratio of unique words to the total number of words across the top $N$ terms from all aspects:
$\text{Diversity}=\frac{\text{Number of unique words among top }N\text{ terms}}{N\times K}$ (10)
A higher Diversity score indicates that each aspect captures distinct and non-overlapping semantic concepts. This suggests that the learned aspect embeddings are well separated in the semantic space.
Embedding Coherence (EC) complements traditional co-occurrence-based metrics by measuring the semantic similarity among top aspect terms in the embedding space. It is defined as the average pairwise cosine similarity among the embeddings of the top $N$ aspect terms:
$\text{EC}(W)=\frac{1}{\binom{N}{2}}\sum_{i=1}^{N-1}\sum_{j=i+1}^{N}\cos(\text{Vec}(w_{i}),\text{Vec}(w_{j}))$ (11)
A higher EC value indicates that the embeddings of words within the same aspect category are semantically closer to each other. Unlike simple co-occurrence statistics, this metric reflects not only the frequency of word co-occurrence but also their semantic orientation and contextual similarity in the embedding space, providing a more comprehensive measure of how coherently the model captures semantic concepts.
To verify the effectiveness of the unsupervised aspect representation learned by MUSCAD, we conducted comparative experiments with nine representative unsupervised aspect extraction methods: LDA, BTM, NMF, BERTopic, Top2Vec, CTM, KeyBERT, SimCSE, and ABAE. Further details and implementation notes for the nine comparative models can be found in D. All models were configured with the same number of aspect categories and aspect terms to ensure a fair comparison. In addition to quantitative evaluation, we also visualized the top keywords extracted by each model and performed qualitative comparative analysis to assess both the interpretability and semantic coherence of the aspects generated by each model.
Table 10 demonstrates that MUSCAD outperforms existing unsupervised learning-based methods across all three domains (Hotel, Food, and Beauty).  MUSCAD achieves the highest scores in NPMI, Diversity, and Embedding Coherence, while also maintaining stable UMass values close to zero. These results indicate that MUSCAD effectively learns aspect representations that are semantically coherent, non-redundant, and contextually well-aligned. In addition, the average rank (Rank Avg) is the lowest across all domains, confirming the overall balance and robustness of the model. These findings show that MUSCAD exhibits higher interpretability than probabilistic models and greater contextual coherence than embedding-based models.
To gain deeper qualitative insights into the differences among models, we compared the top 20 Aspect Terms belonging to the scent Aspect Category in the Beauty domain. As shown in Table 11, traditional probabilistic models such as LDA and BTM tend to generate generic and loosely related words (e.g., smell, scent, fragrance, soap, lotion) rather than domain-specific fragrance terms. Embedding-based approaches such as BERTopic and KeyBERT capture a broader contextual range but often mix semantically irrelevant attributes (e.g., hair, body, red, clean), leading to reduced coherence within the aspect cluster.
In contrast, MUSCAD produces a more semantically cohesive and domain-reflective set of terms, including vanilla, lavender, sandalwood, jasmine, and cedarwood, which capture fine-grained nuances of fragrance-related semantics. These findings demonstrate that MUSCAD can consistently learn coherent and semantically aligned Aspect representations across domains in an unsupervised setting, while generating interpretable and fine-grained Aspect structures that align well with human perception.
Model Aspect Term
LDA smell, scent, perfume, pleasant, fresh, soap, clean, shower, body, wash, creamy, feel, nice, water, mild, chemical, moisturize, absorbed, soft, clear
BTM scent, buy, make, start, beauty, nice, best, careful, lip, girl, tell, show, start, say, turn, let, make, mix, husband, hair
NMF smell, scent, fragrance, perfume, pleasant, fresh, sweet, soap, lotion, oil, soft, clean, creamy, smooth, silky, moisturizing, chemical, body, wash, lingers
CTM lovely, smell, pleasant, scent, sweet, subtle, overpowering, nice, wonderful, nicely, creamy, smoothly, fragrance, strong, floral, medium, gorgeous, blush, lightweight, matte
BERTopic hair, skin, smell, shampoo, well, cream, soft, oil, natural, every, body, red, clean, soap, iron, light, think, leaf, wear, last
Top2Vec scent, fragrance, perfume, smell, soap, lotion, wash, spray, shampoo, fresh, coconut, oily, greasy, moisturizing, moisturizer, clean, shower, creamy, silky, pleasant
KeyBERT odor, smell, perfume, fragrance, greasy, leaf, invigorating, nice, clean, fresh, soap, body, lotion, water, skin, hair, color, fair, powdery, pleasant
SimCSE smell, scent, like, love, product, good, fragrance, nice, great, really, oil, feel, perfume, strong, doesnt, light, dont, pleasant, skin, little
ABAE aroma, mint, fruity, floral, overpowering, scent, vanilla, sweet, smell, musk, flowery, scented, feminine, smelling, pleasing, musky, flavor, masculine, hint, vibe
MUSCAD pungent, fruity, flowery, musky, woodsy, spicy, floral, earthy, minty, sweet, citrusy, patchouli, sandalwood, vanilla, lavender, peppermint, tabacco, bergamot, jasmine, cedarwood
Table 11: Comparison of Top 20 Aspect Terms for Scent Category in Beauty Domain by Different Topic Modeling Methods
@@ 5.8 Summary
As demonstrated through various experiments in Section 5, the multi-aspect labels generated by the proposed framework are of sufficient quality for use in model training. The multi-label classification experiments using various fine-tuned models yielded high Micro-F1 and Macro-F1 scores, indicating that the automatically labeled data is suitable for supervised learning. Furthermore, we compared the performance of fine-tuned models with few-shot labeling using GPT-3.5-turbo and GPT-4o-mini to evaluate the capabilities of LLMs in multi-aspect labeling. The results showed that while LLMs were able to assign appropriate labels in some cases, their overall performance was inferior to that of the fine-tuned models. LLMs also exhibited issues with reproducibility and incurred higher costs. In addition, Human Evaluation confirmed that the automatic labels produced by MUSCAD achieved a level of reliability and consistency comparable to manual labeling by experts, and were more effective in reducing subjectivity among evaluators. These findings suggest that the proposed framework offers a more practical solution for applying multi-aspect labeling to large-scale review data. In addition to the supervised evaluation, we further validated the effectiveness of MUSCAD in learning unsupervised aspect representations. Comparative experiments with nine representative unsupervised aspect extraction methods—LDA, BTM, NMF, CTM, BERTopic, Top2Vec, KeyBERT, SimCSE, and ABAE—demonstrated that MUSCAD consistently achieved the highest scores across multiple coherence and diversity metrics (NPMI, UMass, Diversity, and Embedding Coherence). These results confirm that MUSCAD learns semantically coherent, non-redundant, and domain-reflective aspect representations, outperforming existing probabilistic and embedding-based approaches. Taken together, the results of both the supervised and unsupervised evaluations highlight the robustness and generalizability of MUSCAD as a comprehensive framework for scalable and interpretable multi aspect representation learning.
@@ 6.1 Conclusion
In this study, we proposed MUSCAD , an unsupervised multi-aspect labeling framework designed to automatically assign aspect categories to multilingual and multi-domain review data. Unlike previous studies that rely on supervised learning approaches requiring large amounts of manually annotated data, the proposed framework integrates K-means–based aspect candidate generation, contextual representation learning using Multi-Head Attention, and classification optimization through Max-Margin Loss, enabling multi-aspect labeling without predefined vocabularies. Experimental results showed that fine-tuned classification models trained on the multi-aspect labeling dataset generated by the proposed framework achieved high F1-scores, demonstrating the reliability of the automatically generated labels. Furthermore, compared to large language models (LLMs), MUSCAD produced more consistent and reproducible labeling outcomes, while Human Evaluation confirmed that the quality of its automatically generated labels was comparable to expert annotations. In addition, unsupervised evaluation experiments revealed that MUSCAD effectively learns semantically coherent and well-separated aspect representations, outperforming representative unsupervised aspect extraction methods. Finally, the MUSCAD framework successfully performed large-scale automatic labeling across multiple languages and domains, thereby demonstrating both its practical applicability and scalability in real-world environments.
@@ 6.2 Future Work
This study proposed and evaluated the performance of the unsupervised MUSCAD framework across diverse domains. However, several directions remain for future improvements and extensions. First, we aim to build a fully automated labeling system using AI agents. Currently, partial manual intervention is required for matching aspect categories. In future work, we plan to eliminate manual processes entirely by incorporating AI agents and establishing an automated feedback loop, ultimately enabling fully autonomous labeling.
Second, developing aspect-aware review summarization is another important direction. MUSCAD can automatically label multiple aspects within a sentence and group sentences that share the same aspect, providing a foundation for aspect-level summarization. Furthermore, such aspect-based summaries can be extended to “AI agent systems” that summarize user logs or review histories to build personalized profiles. For example, aspect summaries such as “service”, “food quality”, and “ambience” generated by MUSCAD can be transformed into aspect-based user profiles that reflect user preferences, which can then be utilized for personalized responses in recommendation systems or conversational agents.
Third, the multi-aspect labeled datasets automatically generated by MUSCAD can be leveraged in various future research directions. Beyond serving as input for Aspect-Based Sentiment Analysis (ABSA) models, they can contribute to enhancing contextual understanding and representation learning in knowledge-based sentiment analysis models such as the Knowledge Graph Augmented Network (KGAN) [41]. Moreover, due to MUSCAD’s scalability across multiple languages and specialized domains, it can provide a foundation for new aspect discovery research even in low-resource environments such as healthcare, legal, or manufacturing domains.
These research directions represent a path for MUSCAD to evolve from a simple unsupervised labeling model into an AI agent–driven, personalization-oriented, and aspect-understanding framework. Through this evolution, MUSCAD can serve as a key foundation connecting aspect-level summarization, user profiling, multilingual and domain-adaptive labeling, and future advances in Aspect-Based Sentiment Analysis (ABSA) and knowledge-based sentiment understanding.
Acknowledgements
This work was supported by the Institute of Information & Communications Technology Planning & Evaluation (IITP) grant funded by the Korea government (MSIT) (RS-2024-00337072), the National Research Foundation of Korea (NRF) grant funded by the Korea government (MSIT) (No. RS2024-0040780), and the Institute of Information & Communications Technology Planning & Evaluation (IITP) grant funded by the Korea government (MSIT) (No. RS-2020-II201373, Artificial Intelligence Graduate School Program, Hanyang University). This research was supported by Basic Science Research Program through the National Research Foundation of Korea(NRF) funded by the Ministry of Education(RS-2025-2548326)
@@ Appendix C Multi-Aspect Category Labeling Results
This appendix presents example results automatically labeled by the proposed MUSCAD based on real-world review data. For each domain (hotel, food, and beauty), ten review instances were selected, and the corresponding extracted multi-aspect categories are provided. The results can be found in Tables C.1, C.2, and C.3.
@@
Appendix D Details of Compared Unsupervised Aspect Extraction Methods
To evaluate the effectiveness of MUSCAD, we compare our model with nine representative unsupervised aspect and topic extraction methods, including probabilistic, hybrid, and embedding-based approaches. All baselines are implemented under the same number of aspect categories and aspect terms for fair comparison.
  1. (1)
    LDA (Latent Dirichlet Allocation): the probabilistic topic model that assumes each document is represented as a mixture of latent topics, where each topic is characterized by a distribution over words. It serves as a fundamental baseline for unsupervised topic discovery [44].
  2. (2)
    BTM (Biterm Topic Model): an extension of LDA that models global word co-occurrence patterns (biterms) across the entire corpus rather than within individual documents, improving performance on short texts [45].
  3. (3)
    NMF (Non-negative Matrix Factorization): a deterministic topic extraction approach that factorizes the document-term matrix into two non-negative matrices representing topics and their associated word distributions. It provides interpretable results without probabilistic assumptions [46].
  4. (4)
    CTM (Combined Topic Model): a hybrid topic model integrating contextual embeddings from pre-trained transformers into classical probabilistic topic modeling, enhancing both coherence and interpretability [47].
  5. (5)
    BERTopic (Bidirectional Encoder Representations for Topic Modeling): a transformer-based neural topic model combining BERT embeddings with dimensionality reduction (UMAP), clustering (HDBSCAN), and class-based TF–IDF to generate semantically meaningful topics [48].
  6. (6)
    Top2Vec (Topic2Vec Representation Model): jointly learns topic and document vectors in a shared semantic space using Doc2Vec embeddings, identifying topics as dense clusters of semantically similar documents [49].
  7. (7)
    KeyBERT (Keyword Extraction using BERT): a keyword extraction model leveraging BERT embeddings and cosine similarity between document and candidate word vectors to identify semantically relevant terms [50].
  8. (8)
    SimCSE (Simple Contrastive Sentence Embedding): a sentence embedding model trained via contrastive learning to produce semantically rich sentence representations, applicable to clustering and topic discovery [51].
  9. (9)
    ABAE (Attention-Based Aspect Extraction): an attention-based neural aspect extraction model that learns unsupervised aspect-specific representations from sentences, serving as one of the earliest neural approaches for unsupervised aspect extraction [19].
Appendix E Prompt
This appendix presents the prompts used for multi-aspect labeling with LLMs in the comparative experiment described in Section 5.5. In this study, we designed appropriate prompts to evaluate whether LLMs can assign correct aspect categories to review sentences under zero-shot and few-shot settings. The prompts are shown in Fig. E.1, Fig. E.2, Fig. E.3, Fig. E.4, and Fig. E.5.